OpenStudy (anonymous):

A recipe for 12 corn muffins calls for 1 cup of flour. The number of muffins that can be made varies directly with the amout of flour use. There are 3 1/3 cups of flour available. How many muffins can be made?

OpenStudy (anonymous):

1 cup = 12 12 * 3 = 36. Not sure how to add the 1/3 in.

OpenStudy (anonymous):

3x12 =36 12/3=4 36+4=??

OpenStudy (anonymous):

40.

OpenStudy (anonymous):

yeap !!!!! do you understand why

OpenStudy (anonymous):

Yes, I believe so. Thank you!